OpenStudy (anonymous):

2x^2 / x^2-x-6.... please graph the function including any x intercepts, y intercepts, vertical and horizontal asymptotes, and maximum or minimums

OpenStudy (anonymous):

set the denominator x^2-x-6=0 then factor to find the vertical asymptotes (x-3)(x+2)=0, then x-3 = 0 and x+2=0, so the vertical asymptotes are x = 3 and x = -2

OpenStudy (anonymous):

to find the x-intercepts set the numerator = 0, or 2x^2 = 0, or x = 2, is the only x intercept

OpenStudy (anonymous):

sorry mistake x intercept is 0

OpenStudy (anonymous):

so is y-intercept, so graph crosses at (0,0) the horizontal asympote is a y=2, since the numerator and the denominator both have the same degree (there largest exponent variable is the same) the asymptotoes is at y=2/1 or the leading coefficient of the numerator / leading coefficient of the denominator